Suppose you decide to save 10% of your net monthly income. Find this amount. Monthly Savings = $270.00

If you invest this money in an account that compounds monthly with an APR of 2.5%, how much will you have saved i) after 1 year? ii) after 3 years?
Answer #1

here we have monthly saving is $270

that is 10% of the net monthly income

.

(1+5) - 1 FV = PMT

PMT=270

r=2.5% = 0.025

n=12 for monthly payment

t=1 year

0.025 12 70 + (1 12 FV 0.025

FV = 270 (1 + 0.002083) 12 - 1 0.002083

FV = 270 (1.002083)2 - 1 0.002083

(1.025288 – 1 FV = 270 0.002083

FV = 270 0.025288 0.002083

FV = 270 12.13845

FV = 3277.38402...................... after 1 year

.

.

.

.

(1+5) - 1 FV = PMT

PMT=270

r=2.5% = 0.025

n=12 for monthly payment

t=3 year

0.025 123 12 FV - 20 / (1+ FV = 270 0.025

FV = 270 (1 +0.002083) 36- 0.002083

(1.0778 - 1 FV = 270 0.002083

FV = 270 0.0778 0.002083

FV = 270 37.34402

FV = 10082.8879...................... after 3 years